# and quantum information science

### v1

@@ 6.2 Krylov complexity of evolving operators
In Sec. 6.1, we followed [17] to quantify how quantum states spread dynamically over the Hilbert space. In this section we will discuss a similar approach to the spread of operators [16], which was introduced earlier than Ref. [17]. We started with a description of the spread of states because it is technically simpler to some extent, and the proofs of minimality of the Krylov basis were developed in that context. But the relation between the operator-based methods of this subsection and the state-based methods of the previous section is analogous to the relation between the Heisenberg and Schrödinger approaches to quantum mechanics – either approach may be more convenient depending on the question we are studying.
The idea of defining operator size through an expansion in a natural basis was, to our knowledge, first explored in holographic systems such as the SYK model [143, 144, 145, 146]. This concept was later quantified in [16], whose authors studied operator spread in the Krylov basis, and proposed to quantify operator size in terms of Krylov complexity (the operator analog of (63)). In this framework, we consider the Heisenberg evolution of an initial, "simple" operator $\mathcal{O}(0)$
@@
where in the second step we introduced the Liouvillian super-operator $\mathcal{L}=[H,\,\cdot\,]$. As before, we can expand the operator $\mathcal{O}(t)$ as a power series in $t$ that involves nested commutators of the Hamiltonian with $\mathcal{O}(0)$ or equivalently various powers of the Liouvillian acting on $\mathcal{O}(0)$. As in Sec. 6.1 this series builds the Krylov subspace to which we want to apply the Lanczos procedure. If the Hilbert space is $N$–dimensional, the associated operator space in which we carry out this procedure is $N^{2}$–dimensional because operators have matrix elements $O_{ab}$, with $a,b=1...,N$.
The procedure for extracting the Krylov basis for operators and the associated Lanczos coefficients is the same as described in Sec. 6.1 for states, after vectorizing operators by mapping them into an auxiliary Hilbert space by the GNS (Gelfand-Naimark-Segal) construction (see, e.g., the discussion in [147]). In quantum information theory such a vectorization also appears as the Choi-Jamiolkowski isomorphism, also known as channel-state duality [148, 149]. After vectorizing operators in this way, $\mathcal{O}\to|\mathcal{O})$, we also have to choose an inner product on the auxiliary Hilbert space. The authors of [147] showed that the Krylov complexity defined with respect to the alternative choices behaves differently as a function of time. However, the approach described in Sec. 6.1 suggests that we should pick the inner product that produces a Krylov basis that minimizes the spread of operators. As discussed in [17] the results of [147] imply this minimization is achieved by the Wightman product. Indeed, this was the choice advocated in [16], albeit for other reasons. Thus, we set
@@
$|\mathcal{O}(t))=e^{i\mathcal{L}}|\mathcal{O}(0))=\sum^{\mathcal{K}-1}_{n=0}i^{n}\varphi_{n}(t)|K_{n}).$ (68)
For Hermitian operators the $a_{n}$’s vanish and we derive the Schrödinger equation
@@
Solving it yields the probability $p_{n}(t)=|\varphi_{n}(t)|^{2}$ and the Krylov complexity that generalizes the notion of the operator size is computed by (63). Interestingly, [150] showed that the dimension $\mathcal{K}$ of the Krylov space for operators is bounded by
@@
In the second equality, we introduced the resolutions of the identity in the energy basis and represented the answer in terms of the matrix elements of the operator.
The spread of states and operators can be used as a probe for understanding the dynamcis of quantum systems. Below we will show how to use this tool in a number of different systems. General results can also be obtained. For example, the authors of the pioneering paper [16] performed numerical and analytical studies of Krylov complexity in integrable and chaotic systems, including the SYK model, and found that the growth of Lanczos coefficients $b_{n}$ is at most linear in $n$
@@
They conjectured that theories in which operator growth saturates this bound are chaotic, in the same sense as probed by the OTOC correlators (see Sec. 2 for a definition of these correlators). In particular, they related the coefficient $\alpha$ to the characteristic exponent $\lambda_{K}$ in an exponential growth of the Krylov complexity for chaotic theories, i.e., $\lambda_{K}=2\alpha$. In the case of the SYK model, this exponent coincided with the Lyapunov exponent read from the OTOC, which we label with $\lambda_{\text{OTOC}}$. In other words, they observed that for the SYK model $\lambda_{K}=\lambda_{\text{OTOC}}$. This match will be further discussed in section 6.3.3. However, in many theories, including 2D CFTs, operator return amplitudes are universal and do not readily distinguish between integrable and chaotic models. Therefore the hypothesis of [16] requires refinement; indeed, as we will discuss below in Sec. 6.3.4, the difference between chaos and integrability may be more apparent in the cross-correlations between the Lanczos coefficients and the effects of these correlations on the dynamics.
@@ 6.3.1 Particles on group manifolds
Explicit solutions of the Lanczos algorithm can be obtained when there are dynamical symmetries governing time evolution on the 1D Krylov chain [16, 151, 134, 152]. Since the Krylov space representation depends on both the Hamiltonian and the initial state, this emergent symmetry need not arise from the fundamental symmetries of the model. As we have seen, Lanczos coefficients can be constructed from the survival amplitude. Within the same model, some scenarios may exhibit dynamical symmetries, while others may not. Below, we present examples in which a dynamical symmetry arises from the motion of a particle on a group manifold [134, 17].
Consider a family of Hamiltonians based on the $\mathbb{R}$) group,
$H=\alpha(L_{-1}+L_{1})+\gamma L_{0}+\delta 1\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [L_{0},L_{\pm 1}]=\mp L_{\pm 1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [L_{1},L_{-1}]=2L_{0}\,.$ (73)
Here $\gamma$, $\alpha$, $\delta$ and $h$ are constants that depend on the details of the physical set-up such as the operator dimension, temperature or the initial state as well as the physical, evolving Hamiltonian. Recall, that in the discrete series representation with scaling dimension $h$, $L_{0}\lvert h,n\rangle=(h+n)\lvert h,n\rangle$, $L_{-1}\lvert h,n\rangle=\sqrt{(n+1)(2h+n)}\lvert h,n+1\rangle$, and $L_{+1}\lvert h,n\rangle=\sqrt{n(2h+n-1)}\lvert h,n-1\rangle$. Now suppose we start at $t=0$ with a highest weight state. Then, applying the recursion method or the moment method, the Krylov basis elements are $\lvert K_{n}\rangle=\lvert h,n\rangle$ by construction. Using this we also find that
@@
This form also appears in other contexts such as the SYK model at large $q$. Indeed, for the growth of operator with dimension $h\sim 1/q$ we have $\gamma=0$ and $\alpha=\pi/\beta$. Quantum speed limits have also been used to show that this form leads to the fastest possible universal growth of operators [153]. More generally, (75) can be divided into three different classes depending on the relative magnitude of parameters. The class with $\gamma<2\alpha$ where complexity grows exponentially, the class with $\gamma>2\alpha$ where complexity is periodic and the intermediate class with exactly $\gamma=2\alpha$ where the spread complexity grows quadratically.
For compact semi-simple Lie groups, we generically end up with finite dimensional Krylov spaces. For example, we can consider the SU(2) symmetry and discrete series representations labeled by the spin $j$. There are $12j+$ 1 Lanczos coefficients that read
@@
Again, the constants depend on details of the physical setup. After solving (53) analytically, we find the corresponding complexity is periodic in time
@@
Finally, we can work out a similar class of solutions controlled by the Heisenberg-Weyl symmetry for which the Lanczos coefficients are
@@
As a consequence, for $\delta=\gamma=0$ the evolving state can be recast into the form
$\lvert\psi(t)\rangle=e^{\alpha(a^{\dagger}+a)t}\lvert 0\rangle,$ (80)
and its Krylov complexity grows quadratically in time
$\mathcal{C}_{K}(t)=\alpha^{2}t^{2}.$ (81)
In [151] some of these solutions were found by mapping the Schrödinger equation into the Toda system that was then solved analytically. In [152], the symmetry approach was explained by recognizing that the Schrödinger equation can be effectively written as a three-term recursion relation for orthogonal polynomials in the Liouvillian (or the Hamiltonian). Consequently, when the Lanczos coefficients have a particular dependence on $n$, which makes this equation coincide with the defining relation of known polynomial families, analytical solutions for the wave functions and complexity can be obtained. For example the $\mathbb{R}$) family with $b_{n}$ given by (76) corresponds to Meixner orthogonal polynomials.
@@ 6.3.2 Operator growth in spin chains
Extensive numerical studies of operator growth and Krylov complexity have been conducted in [16, 133, 150, 154, 155]. In generic finite-dimensional, non-integrable systems, the growth of the Lanczos coefficients $b_{n}$ with $n$ exhibits three distinct regimes, which determine the scaling behavior of Krylov complexity. Denoting the number of degrees of freedom by $K$, the first regime is characterized by a linear growth of $b_{n}$’s, leading to the exponential growth of Krylov complexity. For infinite-dimensional systems, this exponential growth persists indefinitely. For finite dimensional systems, after times of order $t\geq\log(K)$ or Krylov index $n\sim K$, the Lanczos coefficients saturate at a constant value $b_{n}\sim\Lambda K$, where $\Lambda$ is a parameter related to the system’s spectral bandwidth. In this regime, Krylov complexity grows linearly with time. Finally, at exponentially large times $t\sim e^{K}$, the Lanczos coefficients gradually decrease toward zero, leading to the saturation of Krylov complexity.
@@
Moreover, the authors of [155] examined how operator growth differs between integrable and chaotic spin-chain models. The authors studied the XXZ spin chain with integrability-breaking deformations, allowing them to interpolate between the two regimes. Noting that the behavior of the Krylov complexity depends both on the Hamiltonian and the choice of initial operator, they focused on examples like the sum of the pair of spin operators located at some fixed distance from the ends of the chain. In these examples they found that the late-time plateau value is suppressed in the integrable phase, whereas in the chaotic phase, it is larger and eventually matches the prediction of random matrix theory for a system of the same dimension. The authors argued that this suppression arose from a sort of Anderson localization in the one dimensional Krylov chain arising from disorder in the Lanczos coefficients. More details can be found in the review [3].
@@
However, at least for generic initial operators/states, the plateau value cannot serve as signature of chaos vs. integrability, for the reasons described below. Recall, from Sec. 6.1.4 that there is an analytical formula relating the density of states and the coarse-grained Lanczos spectrum. This results applied to Hamiltonians and initial states for which the Lanczos coefficients have a continuous large system limit, such as those arising for generic initial states with continuous support in the energy basis. The authors of [156] showed that for such states, the late time plateau of the spread complexity can be determined analytically in terms of the coarse-grained Lanczos spectrum. Since the relation between this spectrum and the density of states is universal, i.e., it applies to both integrable and chaotic theories which can have the same density of states, the plateau value cannot generally separate integrability and chaos, although it may do so for specially prepared initial conditions. That said, as will see in Sec. 6.3.4, the central phenomenon driving the results of [155], i.e., order vs disorder in the Lanczos coefficients, can be distilled into a sharp conjecture regarding the covariances of the Lanczos coefficients for chaotic theories [126, 156]. These covariances affect the approach of the spread complexity to the late time plateau even in cases where the plateau value is the same between integrable and chaotic theories. We will describe this conjecture and compare explicitly with integrable and chaotic spin chains in Sec. 6.3.4

### v2

@@ 6.2 Krylov complexity of evolving operators
In Sec. 6.1, we followed [22] to quantify how quantum states spread dynamically over the Hilbert space. In this section, we will discuss a similar approach to the spread of operators [21], which was introduced earlier than Ref. [22]. We started with a description of the spread of states because it is technically simpler to some extent, and the proofs of minimality of the Krylov basis were developed in that context. But the relation between the operator-based methods of this subsection and the state-based methods of the previous section is analogous to the relation between the Heisenberg and Schrödinger approaches to quantum mechanics – either approach may be more convenient depending on the question we are studying.
The idea of defining operator size through an expansion in a natural basis was, to our knowledge, first explored in holographic systems such as the SYK model [163, 164, 165, 166]. This concept was later quantified in [21], whose authors studied operator spread in the Krylov basis, and proposed to quantify operator size in terms of the Krylov complexity (the operator analog of (73)). In this framework, we consider the Heisenberg evolution of an initial, “simple” operator $\mathcal{O}(0)$
@@
where in the second step we introduced the Liouvillian super-operator $\mathcal{L}=[H,\,\cdot\,]$. As before, we can expand the operator $\mathcal{O}(t)$ as a power series in $t$ that involves nested commutators of the Hamiltonian with $\mathcal{O}(0)$, or equivalently, various powers of the Liouvillian acting on $\mathcal{O}(0)$. As in Sec. 6.1, this series builds the Krylov subspace to which we want to apply the Lanczos procedure. If the Hilbert space is $N$–dimensional, the associated operator space in which we carry out this procedure is $N^{2}$–dimensional because operators have matrix elements $O_{ab}$, with $a,b=1...,N$.
The procedure for extracting the Krylov basis for operators and the associated Lanczos coefficients is the same as described in Sec. 6.1 for states, after vectorizing operators by mapping them into an auxiliary Hilbert space by the GNS (Gelfand-Naimark-Segal) construction (see, e.g., the discussion in [167]). In quantum information theory, such a vectorization also appears as the Choi-Jamiolkowski isomorphism, otherwise known as channel-state duality [168, 169]. After vectorizing operators in this way, $\mathcal{O}\to|\mathcal{O})$, we also have to choose an inner product on the auxiliary Hilbert space. The authors of [167] showed that the Krylov complexity defined with respect to the alternative choices behaves differently as a function of time. However, the approach described in Sec. 6.1 suggests that we should pick the inner product that produces a Krylov basis that minimizes the spread of operators. As discussed in [22], the results of [167] imply that this minimization is achieved by the Wightman inner product. Indeed, this was the choice advocated in [21], albeit for other reasons. Thus, we set
@@
$|\mathcal{O}(t))=e^{i\mathcal{L}t}|\mathcal{O}(0))=\sum^{\mathcal{K}-1}_{n=0}i^{n}\varphi_{n}(t)|K_{n}).$ (78)
For Hermitian operators, the $a_{n}$’s (recall Eq. (63)), vanish and we derive the Schrödinger equation
@@
Solving it yields the probability $p_{n}(t)=|\varphi_{n}(t)|^{2}$ and the Krylov complexity, which quantifies the operator size, is computed by (73). Interestingly, [170] showed that the dimension $\mathcal{K}$ of the Krylov space for operators is bounded by
@@
In the second equality, we introduced resolutions of the identity in the energy basis and represented the answer in terms of the matrix elements of the operator.
The spread of states and operators can be used as a probe for understanding the dynamics of quantum systems. Below, we will show how to use this tool in a number of different systems. General results can also be obtained. For example, the authors of the pioneering paper [21] performed numerical and analytical studies of Krylov complexity in integrable and chaotic systems, including the SYK model, and found that the growth of Lanczos coefficients $b_{n}$ is at most linear in $n$
@@
They conjectured that theories in which operator growth saturates this bound are chaotic, in the same sense as probed by the OTOC correlators (see Sec. 2 for a definition of these correlators). In particular, they related the coefficient $\alpha$ to the characteristic exponent $\lambda_{K}$ in an exponential growth of the Krylov complexity for chaotic theories, i.e., $\lambda_{K}=2\alpha$. In the case of the SYK model, this exponent coincided with the Lyapunov exponent read from the OTOC, which we label with $\lambda_{\text{OTOC}}$. In other words, they observed that for the SYK model $\lambda_{K}=\lambda_{\text{OTOC}}$. This match will be further discussed in section 6.3.3. However, in many theories, including 2D CFTs, operator return amplitudes are universal and do not readily distinguish between integrable and chaotic models. Therefore, the hypothesis of [21] requires refinement; indeed, as we will discuss below in Sec. 6.3.4, the difference between chaos and integrability may be more apparent in the cross-correlations between the Lanczos coefficients and the effects of these correlations on the dynamics.
@@ 6.3.1 Particles on group manifolds
Explicit solutions of the Lanczos algorithm can be obtained when there are dynamical symmetries governing the time evolution on the 1D Krylov chain [21, 22, 171, 153, 172]. Since the Krylov space representation depends on both the Hamiltonian and the initial state, this emergent symmetry need not arise from the fundamental symmetries of the model. As we have seen, the Lanczos coefficients can be constructed from the survival amplitude. Within the same model, some scenarios may exhibit dynamical symmetries, while others may not. Below, we present examples in which a dynamical symmetry arises from the motion of a particle on a group manifold [153, 22].
Consider a family of fictitious Hamiltonians based on the $\mathbb{R}$) group,
$H=\alpha(L_{-1}+L_{1})+\gamma L_{0}+\delta\mathbbm{1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [L_{0},L_{\pm 1}]=\mp L_{\pm 1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ [L_{1},L_{-1}]=2L_{0}\,.$ (83)
Here $\gamma$, $\alpha$, $\delta$ are constants that depend on the details of the physical set-up such as the operator dimension, temperature or the initial state as well as the physical, evolving Hamiltonian. Recall, that in the discrete series representation with scaling dimension $h$, $L_{0}\lvert h,n\rangle=(h+n)\lvert h,n\rangle$, $L_{-1}\lvert h,n\rangle=\sqrt{(n+1)(2h+n)}\lvert h,n+1\rangle$, and $L_{+1}\lvert h,n\rangle=\sqrt{n(2h+n-1)}\lvert h,n-1\rangle$. Now suppose we start at $t=0$ with a highest weight state. Then, applying the recursion method or the moment method, the Krylov basis elements coincide with the Lie algebra basis $\lvert K_{n}\rangle=\lvert h,n\rangle$ by construction. Using this we find that
@@
This form also appears in other contexts such as the SYK model at large $q$. Indeed, for the growth of operators with dimension $h\sim 1/q$, we have $\gamma=0$ and $\alpha=\pi/\beta$ [153]. Quantum speed limits have also been used to show that this form leads to the fastest possible universal growth of operators [173]. More generally, (85) can be divided into three different classes depending on the relative magnitude of the parameters. The class with $\gamma<2\alpha$ where complexity grows exponentially, the class with $\gamma>2\alpha$ where complexity is periodic and the intermediate class with exactly $\gamma=2\alpha$ where the complexity grows quadratically.
For compact semi-simple Lie groups, we generically end up with finite dimensional Krylov spaces. For example, we can consider the SU(2) algebra with Krylov basis representation of a tri-diagonal Hamiltonian
$H=\alpha(J_{+}+J_{-})+\gamma J_{0}+\delta\mathbbm{1}\quad;\quad[J_{0},J_{\pm}]=\pm J_{\pm}\quad;\quad[J_{+},J_{-}]=2J_{0}\,.$ (86)
Evolution of the initial highest (lowest) weight state $\lvert j,-j\rangle$ labeled by spin $j$ allows us to express the dynamics using coherent states of SU(2) and discrete series representations labeled by the spin $j$. Indeed, there are $12j+$ 1 Krylov basis vectors and Lanczos coefficients that read
@@
Again, constants $\gamma$, $\alpha$, $\delta$ and $j$, depend on details of the physical setup, i.e., the initial state, and physical Hamiltonian. After solving (63) analytically, we find the corresponding complexity is periodic in time
@@
Finally, we can work out a similar class of solutions controlled by the Heisenberg-Weyl algebra for which the Hamiltonian in the Krylov basis can be represented as
$H=\alpha(a^{\dagger}+a)+\gamma N+\delta\mathbbm{1}\quad;\quad[a,a^{\dagger}]=1,\quad[N,a^{\dagger}]=a^{\dagger},\quad[N,a]=-a\,,$ (89)
where $N=a^{\dagger}a$. The Hamiltonian is tri-diagonal in the $\lvert n\rangle$ basis of eigenstates of the number operator $N$, which is generated by the ladder operators $a$ and $a^{\dagger}$. Thus, this basis is the infinite-dimensional Krylov basis for $H$. The Lanczos coefficients are
@@
For $\gamma\to 0$ it grows quadratically, but otherwise oscillates in time despite the Krylov basis being infinite-dimensional. Observe also that, in all the three examples, spread complexity is independent of $\delta$.
Let us stress again that the examples and Hamiltonians $H$ that we discussed above are effective descriptions of the state dynamics on the Krylov chain. Namely, there are physical setups where the evolution
$\lvert\psi(t)\rangle=e^{-iH_{phys}t}\lvert\psi_{0}\rangle\,,$ (92)
yields a return amplitude $S(t)=\langle\psi(t)|\psi_{0}\rangle$, from which the moments and Lanczos coefficients can be fit into the forms described above and $\lvert\psi(t)\rangle$ is effectively a coherent state for discrete series representations of some semi-simple Lie algebra. This way, the dynamics can be interpreted as motion on group manifolds governed by the Hamiltonians (83), (86) or (89). Although such analytical scenarios are not generic and we have to rely on numerics, there are several interesting examples that fall into this class and we will discuss some of them below.
In [171] some of these solutions were found by mapping the Schrödinger equation into the Toda system that was then solved analytically. In [172], the symmetry approach was explained by recognizing that the Schrödinger equation can be effectively written as a three-term recursion relation for orthogonal polynomials in the Liouvillian (or the Hamiltonian). Consequently, when the Lanczos coefficients have a particular dependence on $n$, which makes this equation coincide with the defining relation of known polynomial families, analytical solutions for the wave functions and complexity can be obtained. For example, the $\mathbb{R}$) family with $b_{n}$ given by (87) corresponds to Meixner orthogonal polynomials.
@@ 6.3.2 Operator growth in spin chains
Extensive numerical studies of operator growth and Krylov complexity have been conducted in [21, 152, 170, 174, 175]. In generic finite-dimensional, non-integrable systems, the growth of the Lanczos coefficients $b_{n}$ with $n$ exhibits three distinct regimes, which determine the scaling behavior of Krylov complexity. Denoting the number of degrees of freedom by $K$, [^22] In some of the references, the number of degrees of freedom is denoted by $S$ and the size of the Hilbert space by $D$. We traded those symbols for $K$ and $N$ respectively, for consistency with other parts of the review. The only exception is the section about SYK where the number of degrees of freedom is denoted $N$ for consistency with the large literature on the topic. the first regime is characterized by a linear growth of $b_{n}$’s, leading to the exponential growth of Krylov complexity. For infinite-dimensional systems, this exponential growth persists indefinitely. For finite dimensional systems, after times of order $t\geq\log(K)$ or Krylov index $n\sim K$, the Lanczos coefficients saturate at a constant value $b_{n}\sim\Lambda K$, where $\Lambda$ is a parameter related to the system’s spectral bandwidth. In this regime, Krylov complexity grows linearly with time. The above-mentioned time evolution of the Lanczos coefficients resembles a switchback effect, see Fig. 5. Indeed, a linear growth is achieved after an initial time, logarithmic in the number of degrees of freedom. Finally, at exponentially large times $t\sim e^{K}$, the Lanczos coefficients gradually decrease toward zero, leading to the saturation of Krylov complexity.
@@
Moreover, the authors of [175] examined how operator growth differs between integrable and chaotic spin-chain models. The authors studied the XXZ spin chain with integrability-breaking deformations, allowing them to interpolate between the two regimes. Noting that the behavior of Krylov complexity depends on both the Hamiltonian and the choice of initial operator, they focused on examples of initial operators such as the sum of two spin operators located at some fixed distance from the ends of the chain. In these examples, they found that the late-time plateau value is suppressed in the integrable phase, whereas in the chaotic phase, it is larger and eventually matches the prediction of random matrix theory for a system of the same dimension. The authors argued that this suppression arises from a sort of Anderson localization in the one-dimensional Krylov chain arising from disorder in the Lanczos coefficients. More details can be found in the review [3].
@@
However, at least for generic initial operators/states, the plateau value cannot serve as a signature of chaos vs. integrability, for the reasons described below. Recall from Sec. 6.1.4 that there is an analytical formula relating the density of states and the coarse-grained Lanczos spectrum. This formula applies to cases where the Hamiltonian and initial state lead to Lanczos coefficients that have a continuous large system limit, for example, if we are working with generic initial states with continuous support in the energy basis. The authors of [24] showed that for such states, the late time plateau of the spread complexity can be determined analytically in terms of the coarse-grained Lanczos spectrum. Since the relation between this spectrum and the density of states is universal, i.e., it applies to both integrable and chaotic theories which can have the same density of states, the plateau value cannot generally separate integrability and chaos, although it may do so for specially prepared initial conditions. That said, as we will see in Sec. 6.3.4, the central phenomenon driving the results of [175], i.e., order vs disorder in the Lanczos coefficients, can be distilled into a sharp conjecture regarding the covariances of the Lanczos coefficients for chaotic theories [23, 24]. These covariances affect the approach of the spread complexity to the late time plateau even in cases where the plateau value is the same between integrable and chaotic theories. We will describe this conjecture and compare its predictionsexplicitly with results for integrable and chaotic spin chains in Sec. 6.3.4.